OpenStudy (anonymous):

where is 1/(1+cosx) not differentiable and why??

OpenStudy (anonymous):

you can diffrentiate it: the answer is, -sinx/(1+cosx)^2

OpenStudy (nikvist):

\[\cos{x}=-1\quad,\quad x=(2k+1)\pi\quad k\in\mathbb{Z}\]